OpenStudy (anonymous):

#1: You have four fixed-volume containers at STP. Container A has 0.5 mol of gas in 11.2 L. Container B has 2 mol of gas in 22.4 L. Container C has 1 mol of gas in 22.4 L. Container D has 2 mol of gas in 11.2 L. Which of the four containers have equal pressures? A. A and B B. A and C C. B and C D. C and D **my answer: B. A and C is that right? @UnkleRhaukus :)

OpenStudy (anonymous):

#2: Why does opening the air valve of a tire at a constant temperature decrease the pressure? A. The temperature decreases. B. The volume decreases. C. The atmospheric pressure decreases. D. The number of molecules decreases. **my answer: B. The volume decreases. is that right?

OpenStudy (anonymous):

#3: A student increases the temperature of a 200 cm^3 balloon from 60°C to 180°C. What will the new volume of the balloon be? A. 100 cm^3 B. 236 cm^3 C. 272 cm^3 D. 600 cm^3 **my answer: B. 236 cm^3 is that right? :)

OpenStudy (anonymous):

#4: In a reaction between vinegar and antacid tablets, the antacid is the limiting reagent. At constant pressure and temperature, three tablets produce 600 cm^3 of gas. What volume will four tablets produce? A. 300 cm^3 B. 600 cm^3 C. 800 cm^3 D. 1,200 cm^3 **my answer: C. 800 cm^3 is that correct?

OpenStudy (anonymous):

#5: You breathe in 12.0 L of pure oxygen at 298 K and 1,000 kPa to fill your lungs. How many moles of oxygen did you take in? (Use the ideal gas law: PV = nRT where R = 8.31 L-kPa/mol-K.) A. 0.05 moles B. 1.21 moles C. 2.42 moles D. 4.84 moles **my answer; C. 2.42 moles is that right?? @UnkleRhaukus :)

OpenStudy (unklerhaukus):

#1 is right

OpenStudy (anonymous):

:) yay! what about 2-5? :D

OpenStudy (unklerhaukus):

try #2 again

OpenStudy (anonymous):

ohh should it be A. The temperature decreases? and that is why it makes the pressure decrease? :/

OpenStudy (unklerhaukus):

nooope

OpenStudy (unklerhaukus):

i dont think #3 is right either

OpenStudy (anonymous):

aww :( then idk :( and oaky, so for 3 would it be C. 272 cm^3 ??? i wasn't too sure on that one.. :/

OpenStudy (unklerhaukus):

yeah that what i got for #3 because 200*453.15/333.15 =272

OpenStudy (anonymous):

okay! :) and what about #2 then?

OpenStudy (unklerhaukus):

when you open the valvue on the tire, the high pressurised air rushes out

OpenStudy (anonymous):

okay, so that means C. The atmospheric pressure decreases. ???

OpenStudy (unklerhaukus):

nope

OpenStudy (unklerhaukus):

the atmospheric pressure is not going to change because of one tire

OpenStudy (anonymous):

aww :( so basically it's D. The number of molecules decreases. then?? (from process of elimination of my bad reasoning and answers lol) ??

OpenStudy (unklerhaukus):

yeah thats better

OpenStudy (unklerhaukus):

the molecules go out thought the valve

OpenStudy (unklerhaukus):

i think youve got #4 right

OpenStudy (anonymous):

okay! Aweosme :) what about 5?

OpenStudy (unklerhaukus):

\[PV=nRT\\n=\frac {PV}{RT}\]

OpenStudy (anonymous):

so i got #5 wrong? :(

OpenStudy (unklerhaukus):

\[n=\frac{1000[\text{kPa}]\times 12.0 [\text{L}]}{8.3[\text{L kPa /mol K]}298 [\text{K}]}\\=\frac{100\times12}{8.3\times298}[\text{mol}]\]

OpenStudy (anonymous):

okay so 200/2473.4=0.08 ? :/

OpenStudy (unklerhaukus):

ops ** \[\frac{1000\times12}{8.3\times298}\]

OpenStudy (unklerhaukus):

=12000/2473 =

OpenStudy (anonymous):

= 4.85 so the answer is D. 4.84 moles ???

OpenStudy (unklerhaukus):

thats it!

OpenStudy (anonymous):

yay! thanks :)

OpenStudy (anonymous):

yay! thanks :)